OpenStudy (anonymous):

Could someone check my answer Solve the equation: 1/3x+9-2/x+3=2 x=-23/6 x=-3.833 so I thought that x=0,-2

OpenStudy (anonymous):

\[\frac{1}{3x+9}-\frac{2}{x+3}=2\]?

OpenStudy (anonymous):

ya thats one

OpenStudy (anonymous):

you can subtract by writing \[\frac{1}{3x+9}-\frac{2}{x+3}=\frac{1}{3x+9}-\frac{6}{3x+9}=\frac{-5}{3x+9}\] then gett \[\frac{-5}{3x+9}=2\iff -5=2(3x+9)\] and solve that one

OpenStudy (anonymous):

-5=6x+18

OpenStudy (anonymous):

\[-5=6x+18\] \[x=-\frac{23}{6}\] which is what you got

OpenStudy (anonymous):

yes but i don't know if it is x=0,-2 x=0,2,-2 x=2,-2 x=0

OpenStudy (anonymous):

it is none of those maybe the question is different than the one you asked

OpenStudy (anonymous):

those answers suggest that you have to solve a quadratic equation, which this one is not

OpenStudy (anonymous):

oh okay ill try to figure it out